MOLECULAR SYMMETRY AND SPECTROSCOPY: THE SECOND EDITION

\underline{P. R. BUNKER}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario, Canada K1A OR6.; PER JENSEN, FB 9 - Theoretische Chemie, Bergische Universität - Gesamthochschule Wuppertal, D-42097 Wuppertal, Germany..

The first edition of Molecular Symmetry and Spectroscopy by P. R. Bunker was published by Academic Press in 1979. We have now written The Second Edition. It is to be published by The National Research Council of Canada Research Press (http://www.nrc.ca/cisti/journals/mgraphs.html), and it will be available in late 1998. To reflect the very significant progress that has been made in high resolution molecular spectroscopy over the last 20 years it has been necessary to greatly increase the amount of material covered, and the new edition is about 50% longer than the first edition. As before the central theme is the use of symmetry in the understanding of molecules and molecular spectra, with emphasis on the use of the molecular symmetry group. The presentation of the material has been altered so that the molecular symmetry group is introduced early on (in Chapter 3). Many experimental spectra are displayed to illustate the theory. This talk will review the contents of the book.